This is another test. Here we go with some tests. We want to make sure we have examples for all

possibilities. For instance, we believe that xdvi should be able to show a figure that has a bounding box and is included via psfig. Here is an attempt, which we call Fig. 1.

Figure 1: Figure with bounding box included with psfig.

This figure is a crude postscript figure created with an in-house program.

This next figure is created with matlab. Also included courtesy of psfig.

Figure 2: Figure with bounding box included with psfig. More robust postscript.

You will no doubt notice the figure does not seem to be centered. It is in fact centered as you can check by clicking on the “View PS” option and checking out its bounding box. It just so happens that the program that generated this figure comes out with weird bb’s.

Now, we might want to try figures that are somewhat different. For instance Fig. 3 has no bounding box. We wish to include it by means of a special. Here we go...

Figure 3: Figure without a bounding box included with special.

You no doubt noticed that the figure is not centered, either horizontally or vertically. With a special command you would have to do that on your own by specifying the hoffset, voffset, etc. In contrast psfig or epsf will do that for you automagically.

Also, you will notice that since the figure is included with a special, it is impossible to know its bounding box (even if it had one—see next example). Therefore it cannot be shown.

Now the following figure, Fig. 4, does have a bounding box but we still include it with a simple special. Plus it is color!

Figure 4: Figure with a bounding box included with special.

Even though the figure does have a bounding box we cannot show it since it was included with a special which does no filtering of the file whatsoever.

The following figure, Fig. 6, is included with epsf. Figure 5: Postscript code directly from an epsf command.

Figure 6: Postscript code directly from an epsf command (FP).

The following figure, Fig. 7, is included with epsf. It is the same figure as above but a Bounding box is provided.

Figure 7: Postscript code directly from an epsf command.

This example is interesting because it shows that if you are not careful with either the dimensions you think the picture has or the bounding box you use for it (assuming it does not have one), then you may have strange things happen to you.

The following figure, Fig. 8, is an example of raw postscript being sent to the driver. It was taken from the dvips manual.

Figure 8: Postscript code directly from a special command.

It may come as a bit of a surprise, but this was the last type of special command to be implemented. But it has now been implemented. It has no bounding box.